hello1234567890:

Paul runs 4.8 km every morning. How many feet are in 4.8 km, given that 1 mile = 1.609 km and 1 mile = 5,280 feet?

Shadow:

\[\frac{ 4.8km }{ 1 } \times \frac{ 1mile }{ 1.609km } \times \frac{ 5,280 feet }{ 1 mile }\] Multiple across and remember to cancel terms. You're just doing a 3 part conversion.